OpenStudy (anonymous):

Complete the equation of the graphed linear function. Write the slope in decimal form. y = x + The graph of the function goes through point negative 4, negative 1 and point 0, negative 4.

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

Hello @alyssa831 ! :) first let's find the slope \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2) do you know how to do this? :)

OpenStudy (anonymous):

i will try @jigglypuff314

OpenStudy (anonymous):

slope=rise/run=-4,-1/0,-4 right?

jigglypuff314 (jigglypuff314):

not quite :) from (-4, -1) we get x1=-4 and y1=-1 from (0, -4) we get x2=0 and y2=-4 so slope = (-4 - (-1)) / (0 - (-4)) = (-4 + 1) / (0 + 4) = ?

OpenStudy (anonymous):

-3/4 -0.75?

jigglypuff314 (jigglypuff314):

yep! so that's your slope :)

OpenStudy (anonymous):

ok so what next

jigglypuff314 (jigglypuff314):

now we have to find the y-intercept that is when x = 0 I think they already gave that point for your though ;)

OpenStudy (anonymous):

-1 and -4 are the y intercepts i think

jigglypuff314 (jigglypuff314):

hmm only -4 is a y *intercept* because when x is 0, y is -4 the other one is just a y *value* so we have the y intercept = -4 got that? :)

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

then they want you to put those two numbers into the "slope intercept form" which is the y = m x + b where m = slope and b = y intercept can you plug those two numbers in? :)

OpenStudy (anonymous):

so its y=-0.75x+-4

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

thank you

jigglypuff314 (jigglypuff314):

Glad I could help! :)